# White-Basilisk: A Hybrid Model for Code Vulnerability Detection

### v1

@@ 1 Introduction
Current research has explored various approaches to address these challenges. Transformer-based models fine-tuned on code have shown promising results (Feng et al., 2020; Guo et al., 2022), but their quadratic attention complexity limits their application to large codebases. Specialized vulnerability detection architectures (Hanif and Maffeis, 2022; Fu and Tantithamthavorn, 2022) improve upon these foundations but still struggle with long-range dependencies and resource efficiency. Notably, even as these models grow in size and complexity, their performance on realistic vulnerability detection benchmarks often falls short of practical requirements (Ding et al., 2024). This disconnect between model scale and real-world effectiveness suggests an opportunity to re-examine fundamental architectural assumptions in the context of this specialized task.
We introduce White-Basilisk, a compact 200M-parameter model that challenges conventional scaling wisdom by outperforming substantially larger models on code vulnerability detection tasks. Our approach diverges from the standard practice of scaling existing architectures, instead focusing on designing a specialized hybrid architecture that addresses the unique requirements of vulnerability detection. The key insight driving our work is that by combining state-space models (Mamba) with linear-complexity attention mechanisms and sparsely-activated Mixture of Experts, we can efficiently process extremely long code sequences while capturing both local syntactic patterns and global dependencies crucial for vulnerability identification.
@@
    Novel Hybrid Architecture: We propose a resource-efficient model integrating Mamba layers for local pattern recognition, linear-complexity attention for global context modeling, and conditional computation through Mixture of Experts, enabling effective vulnerability detection with just 200M parameters.
@@
    Extended Context Processing: White-Basilisk analyses sequences up to 128,000 tokens during inference—over 8× longer than typical transformer models—enabling comprehensive analysis of entire codebases on a single GPU and detection of vulnerabilities spanning multiple functions or files.
@@
  4. 4.
    Environmental and Computational Efficiency: We provide a detailed analysis of White-Basilisk’s carbon footprint and computational requirements, demonstrating up to 99.9% reduction in emissions compared to larger models while maintaining superior performance—establishing a new efficiency frontier for secure, sustainable AI systems.
Beyond the immediate applications in software security, our work presents broader implications for the field of AI. By demonstrating that a carefully designed domain-specific architecture can outperform general-purpose scaled models, we challenge the notion that parameter count is the primary determinant of model capability. White-Basilisk serves as evidence that the path toward more capable AI systems may lie not only in scaling existing architectures but also in rethinking fundamental design choices for specific high-value applications. In the following sections, we detail White-Basilisk’s architecture, training methodology, and comprehensive evaluation across multiple vulnerability detection benchmarks.
@@ 2 Related Work
Driven by the growing complexity and security risks of software systems, early pioneers explored static analysis techniques Chess and McGraw (2004) and pattern matching methods Livshits and Lam (2005). While these approaches laid a foundational framework, they often encountered significant drawbacks, including high false positive rates, difficulty detecting complex vulnerabilities, and susceptibility to obfuscation techniques. Due to these limitations, they were generally unsuitable for active production environments.
@@
Following that, pre-trained LLMs emerged as a significant breakthrough in code analysis. Hanif and Maffeis (2022) proposed VulBERTa, an adaptation of the RoBERTa model for detecting code vulnerabilities, demonstrating the potential of transfer learning from natural language processing to code analysis. By fine-tuning LLMs pre-trained on extensive corpora of code, this approach quickly gained popularity due to its ability to capture latent patterns in both code structure and semantics. However, similar to many transformer-based models, it suffers from quadratic computational complexity with sequence length, constraining its applicability to large-scale projects.
Another line of research focused on models specifically tailored for code understanding. Feng et al. (2020) introduced CodeBERT, a bimodal pre-trained model for programming language and natural language. This work allowed for a deeper understanding of both code and its associated documentation, proving useful across a range of software engineering tasks, including vulnerability detection. However, while powerful, these models typically demand significant computational resources and encounter difficulties when processing extremely long sequences, which restricts their practical use in analyzing large codebases. For instance, CodeBERT was pre-trained for approximately 1000 hours on 16 interconnected NVIDIA Tesla V100 GPUs, representing a substantial energy and resource investment.
Other recent work concentrated on improving the granularity and efficiency of vulnerability detection. Such an example is LineVul (Fu and Tantithamthavorn (2022)), a transformer-based approach for line-level vulnerability prediction, enabling precise localization of vulnerabilities within codebases. While valuable for precise localization, this approach may overlook vulnerabilities that span multiple lines or functions.
@@
Comprehensive studies by Chakraborty et al. (2021) and Ding et al. (2024) highlighted persistent challenges in the field, including data quality issues, unrealistic evaluation methods, and difficulties in handling long-range dependencies. Ding et al. (2024) showed that existing benchmarks significantly overestimate model performance, with state-of-the-art models achieving high scores on flawed datasets but failing on more realistic ones. Our work directly addresses these concerns through improved data collection, realistic evaluation metrics, and model architecture designed for long-range understanding.
Currently, existing vulnerability detection methods primarily focus on their main objectives without considering the increasing size of model parameters. This uncontrolled growth in model size has led to the development of energy-hungry models requiring dozens of cutting-edge GPUs. Most recently, research developments have been made in efficient language model design. For example, building on the work of Mamba Gu and Dao (2023), Lieber et al. (2024) introduced Jamba, which employs a Mixture of Experts (MoE) strategy, combining Mamba layers and attention mechanisms in an interleaving pattern Fedus et al. (2022). In addition, Wang et al. (2020) proposed the Linformer, which can reduce self-attention complexity to linear time using low-rank approximations, while the Transformer-XL architecture Dai et al. (2019) has been shown to effectively model long-term dependencies by introducing recurrence into self-attention.
@@ 3 Model Architecture
The architecture of White-Basilisk was designed to address three key challenges in code vulnerability detection. Firstly, it tackles the problem of long-range dependencies, as code vulnerabilities often span multiple functions or even files, necessitating a model capable of understanding extensive contexts. Secondly, the architecture strikes a balance between local and global information processing. This dual focus enables both fine-grained understanding of code syntax and broad comprehension of overall program structure, both crucial for effective vulnerability detection. Lastly, the design prioritizes computational efficiency, aiming to create a powerful model that can be deployed in real-world settings without requiring massive computational resources. By addressing these interconnected objectives, our model provides a robust solution for identifying and analyzing potential security flaws in code. To this end, we developed a novel hybrid architecture that combines three main components:
  1. 1.
    Mamba layers: These form the backbone of our model, efficiently capturing local dependencies and providing adaptive computation based on input content.
  2. 2.
    White-Basilisk Self-Attention: Our novel adaptation of linear attention mechanisms enables efficient processing of extremely long sequences, enabling whole-codebase analysis.
  3. 3.
    Mixture of Experts (MoE): This adds dynamic adaptability throughout the network, with MoE layers configured with 8 Experts and 2 Experts per token.
The synergy between these components allows White-Basilisk to process sequences up to 128,000 tokens during inference, a capability that sets it apart in the field of code analysis. This extensive context window enables holistic analysis of entire codebases, potentially uncovering vulnerabilities that span multiple functions or files.
Mamba layers form the backbone of White-Basilisk, chosen for their exceptional efficiency in capturing local dependencies in code sequences. Unlike traditional recurrent neural networks (RNNs) or attention mechanisms, Mamba’s selective state-space mechanism allows for linear-time computation with respect to sequence length. In White-Basilisk, Mamba layers are implemented with a State size ($d_{\text{state}}$) of 16, a Convolution kernel size ($d_{\text{conv}}$) of 4 and an Expansion factor of 2.
The Mixture of Experts (MoE) layers in White-Basilisk introduce dynamic adaptability while maintaining computational efficiency. Configured with 8 Experts and 2 Experts per token, these layers allow the model to activate only a subset of parameters for each input token, significantly reducing computational cost compared to fully-dense models of similar capacity. This sparsity contributes to faster inference times, a critical factor in real-world deployment scenarios.
The layer combination pipeline, inspired by the Jamba model Lieber et al. (2024), allows for a sophisticated interleaving pattern. Specifically, the layer combination is defined by two main configuration parameters determined through experimentation: the attention layer offset (2) and the attention layer period (8). The conjunction of layers in our architecture is defined by the following formula:
$\text{Layer}_{i}=\begin{cases}\text{Attention}(x),&\text{if }(i-2)\bmod 8=0\text{ and }i\geq 2\\ \text{MoE}(x),&\text{if }i\bmod 2=1\\ \text{Mamba}(x),&\text{otherwise}\end{cases}$ (1)
The forward pass through the model can be expressed as $h_{i}$, where $i\in\{0,1,...,L-1\}$ is the layer index and $L$ is the total number of layers in the architecture. Here, $h_{i}$ is the hidden state after the $i$-th layer, and $h_{0}$ is the initial input embedding. The final output of the model $y$ is obtained by applying layer normalization to the last hidden state $h_{L}$:
$h_{i}=\text{Layer}_{i}(h_{i-1})+h_{i-1}\quad\text{and}\quad y=\text{LayerNorm}(h_{L})$ (2)
The residual connections facilitate gradient flow during training and allow for the preservation of information across layers.
3.1 White-Basilisk Self-Attention: Our Novel Linear-Scaling Mechanism
A key innovation in White-Basilisk is our novel self-attention mechanism that achieves linear complexity with respect to sequence length, enabling the model to efficiently process entire codebases up to 128K tokens. Traditional attention mechanisms face a fundamental scaling limitation due to their quadratic complexity, making them computationally prohibitive for analyzing large codebases. Our approach transforms this limitation while maintaining attention quality.
Taking inspiration from recent advances in efficient attention Munkhdalai et al. (2024), we developed a substantially modified implementation that prioritizes holistic sequence understanding rather than segmented processing. The primary distinctions in our approach are:
  1. 1.
    Global Context Accumulation: Unlike segmented approaches that process chunks independently, White-Basilisk Self-Attention accumulates context information across the entire sequence:
    $\text{total}_{\text{mem}}=\sum_{s=1}^{S}A_{\text{mem,s}},\quad\text{total}_{\text{attn}}=\sum_{s=1}^{S}A_{\text{dot,s}}$ (3)
    where $S$ is the total number of segments. This accumulation leads to linear memory growth with sequence length while enabling attention across the full context.
  2. 2.
    Adaptive Global Gating: Our approach implements a novel global gating mechanism that operates on the entire accumulated context:
@@ 3.1 White-Basilisk Self-Attention: Our Novel Linear-Scaling Mechanism
    This allows for adaptive balancing of memory-based and direct attention components across the entire sequence, unlike traditional approaches that gate each segment individually.
The memory-based component of our attention mechanism adaptively compresses and retrieves context information through:
@@
$M_{\text{new}}=M+(\text{ELU}(K)^{T}+1)V,\quad z_{\text{new}}=z+\sum_{i=1}^{L}(\text{ELU}(K_{i})+1)$ (6)
where $M$ is the compressive memory, $z$ is the normalization term, and $L$ is the segment length.
A critical aspect of our implementation is the use of learnable parameters to optimize both the memory accumulation and attention distribution. The memory matrix $M$ incorporates learnable projection matrices that are trained to retain information that might be missed by the local processing of Mamba layers. Through shared residual connections with surrounding Mamba layers (as expressed in the forward pass equations), our attention mechanism effectively learns to complement the local processing strengths of Mamba with global context awareness.
This integration is particularly notable in how it differs from the original Infini-attention approach, which processes segments independently. Our adaptation maintains continuous information flow across the entire sequence through the accumulation of memory states, allowing the model to establish connections between distantly separated code elements. This capability is essential for detecting vulnerabilities that span multiple functions—for example, identifying a buffer allocation in one function and its potential overflow in another much later in the codebase.
4 Experimental Setup
The development and evaluation of source code vulnerability detection models requires a large collection of annotated data samples. In this section, we outline the datasets chosen for this purpose and explain how they were used for both model training and testing purposes. Additionally, we provide a detailed overview of the training methodology used for our model.
4.1 Pretraining
@@
Traditional LLM training methods are generally designed to enable models to comprehend language and its syntax. This is often accomplished through Causal Language Modeling (CLM), where a model learns to predict the next token given its input. Another common approach is based on the Fill in the Middle (FIM) technique, in which random text portions are masked, and the model must reconstruct the missing content. Some advanced source code LLMs combine both methods to increase model flexibility (Li et al., 2023). Similarly, in our work, we employ both techniques during model pre-training on the selected 2M code samples. This process requires approximately 600 hours to complete on a single NVIDIA A100 40GB GPU.
@@
For pretraining data, we initially selected a carefully curated subset of the StarCoder dataset (Li et al., 2023), which includes more than 80 programming languages and consists of 305M files in total. For our study, we focused on C and C++ code samples, using 2M code samples during the pre-training phase. This focused approach allowed us to achieve excellent general performance while maintaining a relatively small pretraining dataset, challenging the notion that massive pretraining datasets are always necessary for high-performance models.
4.2 Fine-tuning and Evaluation
@@
These datasets present significant challenges that validate White-Basilisk’s capabilities. They feature significant class imbalance with the vulnerable class representing only 2.9-9.8% of samples, and diverse vulnerability types covering multiple Common Weakness Enumeration (CWE) categories. To ensure fair comparison with baseline models, we used the identical data splits as reported in the original papers for each dataset.
For fine-tuning, we employed a dataset-specific approach to address the substantial class imbalance present in vulnerability detection. Class weighting was implemented to ensure balanced representation of vulnerable and safe code examples during training. Dataset-specific hyperparameter tuning included learning rates between 1e-5 and 5e-6, with smaller batch sizes (4) for fine-tuning to prevent overfitting.
Given the class imbalance in vulnerability detection, we prioritized metrics that provide a balanced assessment of model performance. F1 Score served as our primary metric, balancing precision and recall. We also evaluated using the Vulnerability Detection Score (VD-S) introduced by (Ding et al., 2024), which measures False Negative Rate. Accuracy, precision, and recall are reported for completeness and comparison with baseline models.
5 Experimental Results: Small Model, Big Impact
To evaluate White-Basilisk’s effectiveness, we conducted extensive experiments across five widely-used vulnerability detection datasets. Tables 1 and 2 present our results compared to state-of-the-art models. White-Basilisk consistently outperforms larger, more resource-intensive models across all benchmarks.
On the challenging PRIMEVUL dataset, White-Basilisk achieved an F1 Score of 0.2907 and VD-S of 0.7239, significantly outperforming larger models like StarCoder2 (7B parameters). Most impressively, on BigVul, our model achieved an F1 Score of 0.9490, accuracy of 0.9942, and VD-S of 0.0398, dramatically surpassing all competitors across all metrics. On the Draper dataset, White-Basilisk set a new standard with an F1 Score of 0.6069, while on REVEAL, it scored an F1 of 0.4934 and accuracy of 0.8988, outperforming previous best results. For VulDeepecker, our model achieved an F1 Score of 0.9388 with the highest precision at 0.9720.
Table 1: BigVul and PRIMEVUL Evaluation Results

### v4

@@ 1 Introduction
Current research has explored various approaches to address these challenges. Transformer-based models fine-tuned on code have shown promising results (Feng et al., 2020; Guo et al., 2022), but their quadratic attention complexity limits their application to large codebases. Specialized vulnerability detection architectures (Hanif & Maffeis, 2022) improve upon these foundations but still struggle with long-range dependencies and resource efficiency. Notably, even as these models grow in size and complexity, their performance on realistic vulnerability detection benchmarks often falls short of practical requirements (Ding et al., 2024). This disconnect between model scale and real-world effectiveness suggests an opportunity to re-examine fundamental architectural assumptions in the context of this specialized task.
We introduce White-Basilisk, a compact 200M-parameter model that challenges conventional scaling wisdom by outperforming substantially larger models on code vulnerability detection tasks. Our approach diverges from the standard practice of scaling existing architectures, instead focusing on designing a specialized hybrid architecture that addresses the unique requirements of vulnerability detection. The key insight driving our work is that by combining state-space models (Mamba) with linear-scaling attention mechanisms and sparsely-activated Mixture of Experts, we can efficiently process extremely long code sequences while capturing both local syntactic patterns and global dependencies crucial for vulnerability identification.
@@
    Novel Hybrid Architecture: We propose a resource-efficient model integrating Mamba layers for local pattern recognition, linear-scaling attention for global context modeling, and conditional computation through Mixture of Experts, enabling effective vulnerability detection with just 200M parameters.
@@
    Efficient Context Processing: White-Basilisk is capable of processing extremely long sequences. It achieves this with a 24-fold reduction in memory consumption when compared to the quadratic attention mechanisms common in standard transformer architectures.
@@
Beyond the immediate applications in software security, our work presents broader implications for the field of AI. By demonstrating that a carefully designed domain-specific architecture can outperform general-purpose scaled models, we challenge the notion that parameter count is the primary determinant of model capability. White-Basilisk serves as evidence that the path toward more capable AI systems may lie not only in scaling existing architectures but also in rethinking fundamental design choices for specific high-value applications.
@@ 2 Related Work
Driven by the growing complexity and security risks of software systems, early pioneers explored static analysis techniques Chess & McGraw (2004) and pattern matching methods Livshits & Lam (2005). While these approaches laid a foundational framework, they often encountered significant drawbacks, including high false positive rates, difficulty detecting complex vulnerabilities, and susceptibility to obfuscation techniques. Due to these limitations, they were generally unsuitable for active production environments.
@@
Following that, pre-trained LLMs emerged as a significant breakthrough in code analysis. Hanif & Maffeis (2022) proposed VulBERTa, an adaptation of the RoBERTa model for detecting code vulnerabilities, demonstrating the potential of transfer learning from natural language processing to code analysis. By fine-tuning LLMs pre-trained on extensive corpora of code, this approach quickly gained popularity due to its ability to capture latent patterns in both code structure and semantics. However, similar to many transformer-based models, it suffers from quadratic computational complexity with sequence length, constraining its applicability to large-scale projects.
@@
Comprehensive studies by Chakraborty et al. (2021) and Ding et al. (2024) highlighted persistent challenges in the field, including data quality issues, unrealistic evaluation methods, and difficulties in handling long-range dependencies. Ding et al. (2024) showed that existing benchmarks significantly overestimate model performance, with state-of-the-art models achieving high scores on flawed datasets but failing on more realistic ones.
To address the computational and architectural challenges identified in existing vulnerability detection approaches, White-Basilisk’s architecture builds upon three established techniques that address complementary challenges in efficient sequence modeling.
State-space models, exemplified by Mamba (Gu & Dao, 2023), offer an alternative to traditional recurrent and attention-based architectures for sequence modeling. Mamba employs a selective state-space mechanism that maintains linear computational complexity with respect to sequence length, contrasting with the quadratic complexity of standard transformer attention. The selective mechanism allows the model to focus on relevant information while efficiently processing long sequences.
Linear attention mechanisms address the computational bottleneck of standard self-attention, which scales quadratically with sequence length. Several approaches have been proposed to achieve linear complexity, including low-rank approximations (Wang et al., 2020) and kernel-based methods. Infini-attention (Munkhdalai et al., 2024) represents a recent advancement in this direction, introducing a memory-based approach that maintains both local attention patterns and compressed representations of distant context. These mechanisms aim to preserve the modeling capabilities of attention while reducing computational requirements, though the effectiveness varies depending on the specific task and sequence characteristics.
Mixture of Experts (MoE) architectures introduce conditional computation by activating only a subset of model parameters for each input token (Fedus et al., 2022). This approach allows models to increase their capacity without proportionally increasing computational cost during inference. Recent implementations like Jamba (Lieber et al., 2024) have demonstrated the effectiveness of combining MoE with both attention and state-space components in a hybrid architecture. The sparse activation pattern in MoE can provide computational efficiency while maintaining model expressiveness.
@@ 3 Model Architecture
Figure 1: White-Basilisk Model architecture
White-Basilisk addresses a fundamental limitation in current language models: the quadratic complexity of standard attention mechanisms. Our architecture introduces a hybrid approach that achieves linear complexity with respect to sequence length while maintaining the representational power necessary for complex reasoning tasks. The core innovation lies in combining three complementary components: Mamba layers, a novel linear-scaling attention mechanism, and Mixture of Experts. This combination enables processing long sequences with substantially reduced memory requirements compared to standard transformers.
3.1 Basilisk Self-Attention
The most significant architectural innovation in White-Basilisk is our linear-scaling attention mechanism, which we term Basilisk Self-Attention. Our approach builds upon recent advances in linear attention, particularly the Infini-attention framework (Munkhdalai et al., 2024), but introduces fundamental modifications that better suit general-purpose language modeling. While the original Infini-attention processes sequences in independent segments, our adaptation maintains global context through cumulative information aggregation across the entire sequence.
The mechanism operates by dividing long input sequences into segments of fixed size $S$ (typically 2,048 tokens). For each segment $s$, we compute both local attention patterns and retrieve information from a compressed global memory. The key insight is that rather than processing segments independently, we accumulate context information across the entire sequence, enabling attention to span arbitrarily long distances while maintaining linear complexity with respect to sequence length.
Mathematically, our attention mechanism can be expressed through the following operations. For a sequence divided into $T$ segments, we maintain running totals of memory-based and attention-based components:
$\text{total}_{\text{mem}}=\sum_{s=1}^{T}A_{\text{mem},s},\quad\text{total}_{\text{attn}}=\sum_{s=1}^{T}A_{\text{dot},s}$ (1)
where $A_{\text{mem},s}$ represents the memory-based attention output for segment $s$, and $A_{\text{dot},s}$ represents the standard dot-product attention output for segment $s$. This accumulation process ensures that information from all previous segments remains accessible to current computations.
The final output combines these accumulated components through an adaptive gating mechanism:
@@ 3.1 Basilisk Self-Attention
where $\beta$ is a learnable parameter that determines the balance between memory-based and direct attention components, and $\odot$ denotes element-wise multiplication.
The memory-based component implements a compressed representation of key-value information from previously processed segments. For each segment, we compute:
@@
where $Q$ represents the query matrix for the current segment, $M$ is the compressive memory matrix, $z$ is a normalization term, $\epsilon$ is a small constant for numerical stability, and ELU is the Exponential Linear Unit activation function. The ELU activation ensures non-negative values, which is crucial for the probabilistic interpretation of attention weights.
The memory matrix $M$ and normalization term $z$ are updated incrementally as each segment is processed:
$M_{\text{new}}=M+(\text{ELU}(K)+1)^{T}V,\quad z_{\text{new}}=z+\sum_{i=1}^{S}(\text{ELU}(K_{i})+1)$ (4)
where $K$ and $V$ are the key and value matrices for the current segment, and $S$ is the segment length.
This approach differs fundamentally from other linear attention methods such as Linformer (Wang et al., 2020), which reduces attention complexity through low-rank approximations, or Performer (Choromanski et al., 2022), which uses random feature maps. While these methods achieve linear complexity, they typically sacrifice some representational capacity. Our memory-based approach maintains rich contextual information through the compressed memory mechanism while achieving $O(nSd+2nd^{2})$ computational complexity, identical to the original Infini-attention. However, our implementation differs in memory usage: whereas the original Infini-attention maintains constant $O(d^{2})$ memory footprint by processing segments in streaming fashion, our approach accumulates segment outputs before final concatenation, resulting in $O(nd)+O(d^{2})$ memory complexity. This design choice trades the bounded memory property for simplified implementation and improved global context integration, though it limits processing to sequences that fit within available memory rather than enabling truly infinite sequence processing. Detailed computational and memory complexity analysis is provided in Appendix F.
3.2 Integration with Mamba and Mixture of Experts
White-Basilisk employs a structured interleaving of three layer types: Mamba layers, our linear-scaling attention layers, and Mixture of Experts (MoE) layers. Mamba layers provide the backbone for local pattern recognition and efficient sequence processing. These layers implement selective state-space modeling with linear computational complexity, making them well-suited for capturing local dependencies in long sequences. The MoE layers introduce adaptive computation by activating only a subset of parameters for each token. This design reduces computational cost compared to dense layers of equivalent capacity while maintaining model expressiveness.
The layer combination follows a systematic pattern inspired by Lieber et al. (2024) and defined by two parameters: attention layer offset ($\alpha=2$) and attention layer period ($\pi=8$). The layer type for position $i$ is determined by:
$\text{Layer}_{i}=\begin{cases}\text{Attention}(x),&\text{if }(i-\alpha)\bmod\pi=0\text{ and }i\geq\alpha\\ \text{MoE}(x),&\text{if }i\bmod 2=1\\ \text{Mamba}(x),&\text{otherwise}\end{cases}$ (5)
This interleaving pattern ensures regular insertion of global attention layers while maintaining efficient local processing through Mamba layers. The MoE layers are interspersed to provide adaptive computation where beneficial.
The forward pass through White-Basilisk can be described as a sequence of transformations applied to the hidden state. Let $h_{i}$ denote the hidden state after layer $i$, where $h_{0}$ represents the input embeddings. Each layer applies its transformation while preserving information flow through residual connections:
$h_{i}=\text{Layer}_{i}(h_{i-1})+h_{i-1}$ (6)
The residual connections serve two crucial purposes: they facilitate gradient flow during training, preventing the vanishing gradient problem in deep networks, and they preserve information across layers, allowing the model to selectively refine representations without losing previously computed features.
The resulting architecture maintains linear complexity with respect to sequence length while providing the representational capacity necessary for complex reasoning tasks. This design enables processing of sequences that would be computationally infeasible with standard transformer architectures, opening new possibilities for applications requiring extensive context understanding.
4 Pretraining
@@
Traditional LLM training methods are generally designed to enable models to comprehend language and its syntax. This is often accomplished through Causal Language Modeling (CLM), where a model learns to predict the next token given its input. Another common approach is based on the Fill in the Middle (FIM) technique, in which random text portions are masked, and the model must reconstruct the missing content. Some advanced source code LLMs combine both methods to increase model flexibility (Li et al., 2023). Similarly, in our work, we employ both techniques during model pre-training on the selected 2M code samples. This process requires approximately 600 hours to complete on 1x NVIDIA A100 40GB GPU and emitted approximately 85 kg of CO2.
@@
For pretraining data, we initially selected a carefully curated subset of the StarCoder dataset (Li et al., 2023), which includes more than 80 programming languages and consists of 305M files in total. For our study, we focused on C and C++ code samples, using 2M code samples during the pre-training phase. Crucially, our pretraining stage used exclusively general-purpose code from the StarCoder dataset with no exposure to any vulnerability detection datasets or vulnerability-specific annotations. This separation ensures that our model’s vulnerability detection capabilities emerge solely from the fine-tuning stage rather than from pretraining exposure to security-related patterns.
5 Fine-tuning and Evaluation
@@
These datasets present significant challenges that validate White-Basilisk’s capabilities. They feature significant class imbalance with the vulnerable class representing only 2.9-9.8% of samples, and diverse vulnerability types covering multiple Common Weakness Enumeration (CWE) categories.
Our fine-tuning methodology follows a rigorous evaluation protocol to prevent overfitting to evaluation characteristics. For each dataset, we fine-tuned our pretrained model on the binary classification task (vulnerable vs. non-vulnerable code) using only the publicly available training splits. During training, we monitored performance on the test splits to select the best-performing checkpoint based on F1 score. Finally, we evaluated the selected checkpoint on the validation splits to obtain our reported results. This approach ensures that our final performance metrics represent genuine held-out evaluation without contamination from validation set characteristics.
To address the substantial class imbalance present in vulnerability detection datasets, we implemented three key techniques and selected appropriate evaluation metrics. For training, we applied class weighting using inverse frequency weighting to balance loss contributions between vulnerable and non-vulnerable samples, employed a Weighted Random Sampler without replacement to ensure balanced representation within each training batch, and incorporated Scale-Invariant Fine-Tuning (SIFT) to improve model robustness against small perturbations. For evaluation, we prioritized metrics that provide balanced assessment of model performance, using F1 Score as our primary metric to balance precision and recall, and the Vulnerability Detection Score (VD-S) introduced by (Ding et al., 2024), which measures False Negative Rate. Detailed descriptions of these techniques are provided in Appendix B.5.
6 Experimental Results
To comprehensively evaluate White-Basilisk’s effectiveness, we conducted experiments across multiple dimensions: dataset-specific performance, cross-dataset generalization, paired vulnerability detection, context length analysis, and vulnerability type classification.
6.1 Main Results: Dataset-Specific and Unified Training
We evaluated White-Basilisk using two training strategies: dataset-specific models (trained and evaluated on individual datasets independently) and a unified model (trained on combined train data spits from all datasets and evaluated on the validation set of each dataset).
Table 1: Performance comparison on PRIMEVUL and BigVul datasets